OpenStudy (anonymous):

∫[(x^4-9)^(-1/3)[4x^3],]dx=

OpenStudy (anonymous):

let \[u=x^4-9\]

OpenStudy (anonymous):

then \[du=4x^3 dx\]

OpenStudy (anonymous):

then \[du=4x^3 dx\]

OpenStudy (anonymous):

\[\int u^{-1/3} du\]

OpenStudy (anonymous):

\[=\frac{3}{2} u^{2/3} +C\]

OpenStudy (anonymous):

substitute back in to get in terms of x

OpenStudy (anonymous):

\[\frac{3}{2} (x^4-9)^{2/3} +C\]

OpenStudy (anonymous):

compute the derivative to check the answer